There are two small valves in the bottom of a tub filled with water. If opened, one valve will empty the tub in three hours; the other will empty it in six hours. If both valves are opened at the same time, how long will it take to empty the tub? Show your work in Algebraic form.
Answer by jorel1380(3719)

1/3+1/6=1/t
2/6+1/6=1/t
3/6=1/2=1/t
t=2
If both valves are opened, the tub will empty in 2 hours..